An object (m=0.5kg) is moving along x-axis with v0 = 1.6m/s. At t =0s, at which object is at thr origin, a force directed along the negative x-axis F = -kt^3 starts to apply to it. At t=2s, the object passes the origin again. Find k ( a constant).


1
Expert's answer
2021-09-16T10:25:39-0400

We can use the relation between force and acceleration as:

"F=m{a}=m\cfrac{dv}{dt} \implies dv=\cfrac{F}{m}dt"


Then, we can integrate by using the limits vi=1.6 m/s at ti = 0s and we will consider that vf= 0 m/s when tf = 2 s:

"\int_{v_i}^{v_f} dv=\int_{t_i}^{t_f} \cfrac{F}{m}\,dt=\int_{t_i}^{t_f} \cfrac{-kt^3}{m}\,dt
\\ {v_f}-{v_i} =-\cfrac{k}{m}\int_{t_i}^{t_f} t^3\,dt=-\cfrac{k}{4m}[t^4_f-t^4_i]
\\ \text{ }
\\ \implies k=\dfrac{4m({v_i}-{v_f})}{t^4_f-t^4_i}"

After substitution, we can go ahead and find the value for k:


"k=\dfrac{4(0.5\,kg)({1.6\frac{m}{s}}-0\frac{m}{s})}{(2\,s)^4-(0\,s)^4}
\\ k=\dfrac{(2)({1.6\frac{kg \cdot m}{s}})}{16\,s^4}
\\\implies k=0.2\cfrac{kg \cdot m}{s^5}=0.2\cfrac{N}{s^3}"


In conclusion, we find that k = 0.2 N/s3.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.